Question 4

Without using finance calculator,

Refer to Johan's investment account, i.e. the monthly deposits made into his investment account:


Year Monthly deposits

2017 SR 500

2018 SR 400

2019 SR 700


Find the value of his investment at the end of 2019 if the investment rate was 12% compounded monthly. Find also the interest earned
1
Expert's answer
2020-05-27T09:28:07-0400

These are annuities. We find by the formula:


"FVA=A\times\frac{(1+r)^n-1}{r}"


"r=\frac{0.12}{12}=0.01"


Find the future value and interests

2017-2019:

"FVA=A\times\frac{(1+r)^n-1}{r}=500\times\frac{(1+0.01)^{36}-1}{0.01}=21538.44"

interests:

"21538.44-500\times36=21538.44-18000=3538.44"


2018-2019

"FVA=A\times\frac{(1+r)^n-1}{r}=400\times\frac{(1+0.01)^{24}-1}{0.01}=10789.39"


interests:

"10789.39-400\times24=10789.39-9600=1189.39"


2019:

"FVA=A\times\frac{(1+r)^n-1}{r}=700\times\frac{(1+0.01)^{12}-1}{0.01}=8877.75"


"8877.75-700\times12=8877.75-8400=477.75"